An LPG tank contains 11.0 kg of butane gas, (C4H10), in a 25.0 L gas tank. What must be the pressure of the gas inside the tank when the temperature is 25 DEGREES CELCIUS?


1
Expert's answer
2022-04-14T04:20:04-0400

PV = nRT

P = nRT / V

n = m/M

P = mRT / MV

P = (11000 * 8.314 * (25 + 273)) / (58 * 0.025) = 18 795 374 Pa = 185.5 atm